C5H12+8O2 ----> 5CO2+6H20

How many grams of CO2 will be produced I'd 2.67 grand of O2 reacts?
1
Expert's answer
2020-11-04T14:24:58-0500

"m_{(CO_2)}=2.67g_{(O_2)}\times\frac{1mol_{(O_2)}}{32.00g_{(O_2)}}\times\frac{5mol_{(CO_2)}}{8mol_{(O_2)}}\times\frac{44.01g_{(CO_2)}}{1mol_{(CO_2)}}=2.30g"


Answer: 2.30 g